OpenStudy (anonymous):

describe the vertical asymptotes and holes for the graph of y=(x-2)/[(x+2)(x-2)]

myininaya (myininaya):

there is a hole at x=2 and x=-2 there is a vertical asymp.

OpenStudy (anonymous):

now what do i do to figure that out?

OpenStudy (anonymous):

Is it \[\frac{x-2}{(x+2)(x-2)} = \frac{1}{x+2} ?\]

myininaya (myininaya):

notice the x-2 cancel x-2=0 implies x cannot 2 since x+2 did not cancel x+2=0 means x=-2 is a veritcal asymp.

OpenStudy (anonymous):

oh ok and how is it a vertical asymp.

myininaya (myininaya):

plug in numbers close to 2 on the left and numbers close 2 on the right to see why

myininaya (myininaya):

oops -2

myininaya (myininaya):

as those numbers get close to -2 the bottom is getting closer to 0 which means it is approaching either negative infinity or positive infinity